# Fusion Transformer with Object Mask Guidance for Image Forgery Analysis

### v1

@@ Abstract
In this work, we introduce OMG-Fuser, a fusion transformer-based network designed to extract information from various forensic signals to enable robust image forgery detection and localization. Our approach can operate with an arbitrary number of forensic signals and leverages object information for their analysis – unlike previous methods that rely on fusion schemes with few signals and often disregard image semantics. To this end, we design a forensic signal stream composed of a transformer guided by an object attention mechanism, associating patches that depict the same objects. In that way, we incorporate object-level information from the image. Each forensic signal is processed by a different stream that adapts to its peculiarities. Subsequently, a token fusion transformer efficiently aggregates the outputs of an arbitrary number of network streams and generates a fused representation for each image patch. These representations are finally processed by a long-range dependencies transformer that captures the intrinsic relations between the image patches. We assess two fusion variants on top of the proposed approach: (i) score-level fusion that fuses the outputs of multiple image forensics algorithms and (ii) feature-level fusion that fuses low-level forensic traces directly. Both variants exceed state-of-the-art performance on seven datasets for image forgery detection and localization, with a relative average improvement of 12.1% and 20.4% in terms of F1. Our network demonstrates robustness against traditional and novel forgery attacks and can be expanded with new signals without training from scratch.
@@ 1 Introduction
To this end, we propose the Object Mask-Guided Fusion Transformer (OMG-Fuser), capable of capturing image forensic traces from an arbitrary number of input signals by leveraging the image semantics and fusing them in multiple processing stages into a single robust forgery detection and localization output. Our fusion network comprises two main modules: the Forensic Signal Streams and the Token-Fusion Module. In the former, every input signal is propagated through a different stream to capture its unique traces. Our main novelty in this stage includes the Object Guided Attention mechanism that exploits external instance segmentation masks to drive the extraction process to attend only to image regions that depict the same objects so as to generate comprehensive object-level representations. Any pretrained instance segmentation model can be employed [23], and we utilize a recently proposed class-agnostic model [36]. The latter module is responsible for the fusion of any arbitrary number of forensic signals. The Token-Fusion Transformer (TFT) combines the various representations of an image region generated by the forensic signal streams into a single representation via several transformer blocks. The fused outputs are analysed by the Long-range Dependencies Transformer (LDT) that captures the relations between the representations of the different image regions. For training the proposed network, we propose the Stream Drop augmentation. This randomly discards some network streams during training to prevent the network from over-attending on specific forensic signals. Our proposed approach is employed both for feature and score level fusion, combining RGB information with two and five forensic signals, respectively. We demonstrate its effectiveness on five popular datasets, compared with several handcrafted and deep learning state-of-the-art approaches. Moreover, we demonstrate its robustness against common perturbations and recent neural filters. An overview of the approach is presented in Fig. 1. We will release the code upon acceptance.
@@ 3.5 Training process
Feature Fusion |
SPAN [30] | 21.4 | 83.3 | 72.5 | 98.1 | 42.5 | 86.5 | 0.4 | 56.2 | 0.2 | 52.8 | 20.8 | 75.0 | 7.5 | 61.5 | 23.6 | 73.3 |
@@
OMG-Fuser (Ours) | 84.5 | 97.2 | 86.1 | 97.2 | 63.1 | 91.4 | 34.2 | 79.1 | 72.2 | 92.1 | 53.9 | 88.3 | 77.0 | 95.4 | 67.3 | 91.5 | |
Score Fus. |
DST-Fusion [21] | 75.3 | 94.8 | 85.4 | 93.5 | 48.8 | 79.2 | 14.7 | 56.9 | 20.4 | 50.7 | 38.6 | 80.9 | 24.9 | 72.8 | 44.0 | 75.6 |
@@
Feature Fusion |
PSCC-Net [47] | 54.0 | 86.9 | 26.6 | 80.8 | 29.7 | 65.7 | 63.6 | 62.0 | 52.9 | 77.8 | 45.4 | 74.6 |
@@
OMG-Fuser (Ours) | 90.7 | 96.5 | 99.2 | 99.9 | 78.8 | 83.5 | 71.4 | 78.2 | 72.0 | 81.7 | 82.4 | 88.0 | |
Score Fus. |
DST-Fusion [21] | 80.3 | 93.83 | 92.3 | 97.9 | 69.0 | 78.4 | 44.4 | 75.1 | 66.0 | 78.5 | 70.4 | 84.7 |
@@ 4.1 Experimental setup
OMG-Fuser provides a network for combining both the score-level outputs of multiple image forensics algorithms as well as multiple low-level image forensic signals. Thus, we experiment with two different variants of our architecture: (i) one that incorporates five recent forgery localization algorithms [30, 74, 12, 42, 25] (OMG-Fuser), and (ii) one that directly fuses two recently proposed learnable forensic cues, namely the DCT stream [42] and the Noiseprint++[25] at the feature-level (OMG-Fuser). We train both variants on 40k forged samples and 32k authentic samples compiled from the datasets provided in [41, 54, 13], while we evaluate them on CASIAv1+ [12], Columbia [29], Coverage [71], NIST16 [24], OpenForensics [43], CocoGlide [25] and DID [73] datasets. We employ six recent competitor methods with publicly available implementations for the feature fusion category [30, 47, 74, 12, 42, 25]. Regarding score-level fusion, we reimplement using the same input signals as OMG-Fuser, two popular methods that support the fusion of multiple signals [7, 21], along with a baseline average fusion approach. We use the SAM [36] pretrained model for instance segmentation. DINOv2 [59] model is used as the vision backbone, which is fine-tuned on our OMG-Fuser variant, but kept frozen on OMG-Fuser. Moreover, we employ the F1-score with a threshold of as the main indicator of performance when deployed in the wild and the AUC as an auxiliary threshold-agnostic metric. More information regarding implementation details, datasets, and algorithms, along with localization results in terms of the F1 metric using the best threshold per image, is provided in the supplementary material.
@@ 4.2 Comparison with the state-of-the-art
Image forgery localization evaluation results are presented in Tab. 1. The reported F1 and AUC metrics are computed on the pixel-level. We see that our implementations outperform all state-of-the-art methods in score- and feature-level fusion. Moreover, in the case of score-level fusion, we show that previous approaches, either based on some statistical frameworks or learning approaches without semantic information, perform poorly on the recent deep learning-based forensic signals, even worse than averaging the signals. This effectively shows that our network is capable of effectively fusing multiple forensic signals by exploiting the object-level information of the image, and at the same time, it is capable of handling low-level forensic traces. Moreover, the high performance in feature-level fusion enables future works on new forensic signals to use our network to decrease the burden of developing purpose-built fusion architectures.
Image forgery detection evaluation results are presented in Tab. 2. Following [25], for methods that do not output an image-level forgery detection score, we compute it as the max of the corresponding forgery localization mask. Also, we omit datasets that contain only forged samples. According to the results, both OMG-Fuser implementations outperform all state-of-the-art methods by a significant margin.
@@
OMG-Fuser | 70.4 | 93.5 | 83.2 | 89.5 | |
Signals |
w/o SPAN | 69.2 | 91.5 | 82.3 | 89.1 |
@@
w/o Stream Drop | 66.5 | 90.5 | 81.1 | 87.3 |
@@ 4.3 Ablation Study
In order to better understand the contribution of each of the major components of our architecture, we performed several ablation studies in two directions. First, we evaluated the contribution of each of the fused forensic signals by training our network from scratch and removing the inputs of the network one at a time. Similarly, we proceeded by evaluating the contribution of each of the key components of our architecture by: replacing the OGA with the standard unmasked self-attention [68], replacing the TFT with an average pooling layer, removing the LDT and Stream Drop components. To better study the fusion capabilities of the architecture, we performed the ablation on the OMG-Fuser, which fuses the output of several forensic algorithms. The results of the ablation study are presented in Tab. 3. Removing a single input signal affects the performance of the network adversely. Thus, the network is capable of learning to exploit even minor additional information introduced by new input streams. Moreover, we see that removing any key component of the architecture considerably impacts performance, highlighting that they are of crucial importance for effective signal fusion.
@@ 4.4 Robustness against Perturbations and Filters
4.5 Stream Expansion
The design of the TFT allows the expansion of an already trained model with additional streams. In Tab. 4, we report the overall performance across all the evaluation datasets when expanding our architecture in three different ways. Starting from a variant of OMG-Fuser trained without the TruFor stream, we expand it i) by training only the additional stream and keeping the rest of the network frozen, ii) by fine-tuning the network for 15% of the initial training epochs, and iii) by training the network from scratch. We see that in the first two cases, the network achieves a performance very close to third by requiring only a portion of the computational resources. This suggests that our model has learned to fuse various forensic signals effectively; therefore, it can be expanded with new signals without training from scratch.
@@ 5 Conclusions
6 Acknowledgements
This work was partially supported by the European Union under the framework of the Horizon Europe vera.ai project (contract No.101070093). Also, the computational resources for this work were partially provided by the MeluXina supercomputer under the EuroHPC program (project: EHPC-DEV-2023D03-008).
@@ 6 Acknowledgements
7 Experimental Setup
7.1 Implementation Details
@@
7.2 Datasets
To train our OMG-Fuser models, we utilize data from three publicly available datasets. We randomly sample 25k forged and 25k authentic images from the synthetic dataset used by CAT-Net [41] due to the big variability of its samples regarding compression qualities and depicted topics. We enrich this set with another 10k in-painted images sampled from the DEFACTO [54] dataset. Also, we include all images from the CASIAv2 [13] dataset into our training set to compensate for low-resolution and low-quality images. We utilize 90% of these data for the actual training of the model and the rest 10% for validation purposes. Despite some datasets consisting of more samples, we observed in our experiments that a further increase in the amount of training data did not yield any significant performance increase. Thus, our architecture requires significantly less training data than the previous state-of-the-art ones [25, 42]. Instead, as highlighted by [25], introducing more variability into the training data was more beneficial. For evaluation, following [30, 47, 25], we have selected five popular datasets, namely the CASIAv1+ [12], COLUMBIA [29], COVERAGE [71], NIST16 [24], OpenForensics [43] datasets, including common cases of image forgery. In contrast to previous works [30, 47], we benchmark our methods on the entire evaluation dataset without making any further assumptions about the type of forgery or utilizing subsets of them. Furthermore, to take into account deep-learning based inpainting operations, we further employed the recently introduced CocoGlide [25] dataset and the deep-learning based inpaintings of the DiverseInpaintingDataset (DID) [73]. A summary of all the datasets used in our research is presented in table 5.
7.3 Image Forensics Algorithms
In our evaluation, we consider only methods with publicly available code in order to fairly evaluate them under the same setup. We employ as inputs to our score-level fusion network, OMG-Fuser, the SPAN [30], MVSS-Net++ [12], CATNetv2 [42], TruFor [25] and IFOSN [74] algorithms, due to their competitive performance and their complementarity with respect to the type of artifacts they detect, i.e., artifacts in the RGB domain, edge artifacts, compression artifacts, noise artifacts, and robustness against online sharing operations, respectively. Moreover, we utilize the Noiseprint++ [25] and the DCT-domain stream [42] as inputs to our feature-level fusion network, OMG-Fuser, two learnable forensic signals that captures noise-related and compression-related artifacts, respectively. For all the aforementioned methods, we utilize the official code implementations provided by the original authors.
@@ 7.3 Image Forensics Algorithms
Feature Fusion |
PSCC-Net [47] | 83.4 | 86.7 | 69.3 | 50.2 | 51.3 | 84.1 | 58.5 | 70.7 |
@@
OMG-Fuser (Ours) | 92.0 | 94.6 | 88.3 | 82.1 | 82.0 | 87.7 | 78.9 | 80.1 | |
Score Fus. |
DST-Fusion [21] | 89.3 | 91.8 | 76.6 | 56.0 | 33.3 | 86.5 | 64.8 | 77.4 |
@@
OMG-Fuser (Ours) | 92.2 | 96.6 | 89.3 | 83.7 | 85.2 | 86.2 | 86.1 | 86.8 |
For comparison with other score-level fusion approaches, we employ the OW-Fusion [7], a deep learning-based fusion approach, and implement it with the same input signals used on OMG-Fuser. Furthermore, in order to take into account the previous statistical fusion frameworks, we reimplement a DST-based fusion framework [21, 61] in Python, using again the same inputs with our score-level fusion implementation. Finally, we use the average of all input signals as a baseline approach.
8 Additional Experiments
@@
Feature-fusion Ablation Studies: In order to examine the contribution of each signal to the feature-level fusion approach, we extended the ablation studies of the main paper to the inputs of the OMG-Fuser. The results are being presented in Tab. 7. They show that the removal of any of the fused inputs degrade the overall performance. Thus, they highlight the ability of our architecture to efficiently extract and combine the complementary information from each input stream.
@@
OMG-Fuser (Ours) | 32.3 ms | |
Score Fusion |
SPAN [30] | 1.34s |
@@
9 Explainability
To better understand which parts of the input signals contribute the most to the fused tokens (eq. 5), we employ the Grad-CAM [65] method. In particular, we compute the gradients of the fused tokens with respect to the different inputs to the TFT in (eq. 4) in order to isolate the token fusion process and determine from which tokens the information propagates to the next stages. We compute the gradients based on the output of the TFT, using the squared -norm of the . In these experiments, we employ the OMG-Fuser variant of our architecture due to the easier interpretation of the input signals. The explainability maps for three samples are presented in Fig. 7. Our architecture has learned to attend to the correctly estimated regions in the input signals based on the ground truth while ignoring the regions of the input signals containing erroneous predictions. Also, our network focuses on the signals that better capture the forged and the authentic regions separately, e.g. on MVSSNet++ for the detection of authentic regions, while exploiting information from the image to resolve ambiguity in the input signals.
10 Discussion on Research Ethics

### v2

@@ Abstract
In this work, we introduce OMG-Fuser, a fusion transformer-based network designed to extract information from various forensic signals to enable robust image forgery detection and localization. Our approach can operate with an arbitrary number of forensic signals and leverages object information for their analysis – unlike previous methods that rely on fusion schemes with few signals and often disregard image semantics. To this end, we design a forensic signal stream composed of a transformer guided by an object attention mechanism, associating patches that depict the same objects. In that way, we incorporate object-level information from the image. Each forensic signal is processed by a different stream that adapts to its peculiarities. A token fusion transformer efficiently aggregates the outputs of an arbitrary number of network streams and generates a fused representation for each image patch. We assess two fusion variants on top of the proposed approach: (i) score-level fusion that fuses the outputs of multiple image forensics algorithms and (ii) feature-level fusion that fuses low-level forensic traces directly. Both variants exceed state-of-the-art performance on seven datasets for image forgery detection and localization, with a relative average improvement of 12.1% and 20.4% in terms of F1. Our model is robust against traditional and novel forgery attacks and can be expanded with new signals without training from scratch. Our code is publicly available at: https://github.com/mever-team/omgfuser
@@ 1 Introduction
To this end, we propose the Object Mask-Guided Fusion Transformer (OMG-Fuser), capable of capturing image forensic traces from an arbitrary number of input signals by leveraging the image semantics and fusing them in multiple processing stages into a single robust forgery detection and localization output. Our fusion network comprises two main modules: the Forensic Signal Streams and the Token-Fusion Module. In the former, every input signal is propagated through a different stream to capture its unique traces. Our main novelty in this stage includes the Object Guided Attention mechanism that exploits external instance segmentation masks to drive the extraction process to attend only to image regions that depict the same objects so as to generate comprehensive object-level representations. Any pretrained instance segmentation model can be employed [23], and we utilize a recently proposed class-agnostic model [36]. The latter module is responsible for the fusion of any arbitrary number of forensic signals. The Token-Fusion Transformer (TFT) combines the various representations of an image region generated by the forensic signal streams into a single representation via several transformer blocks. The fused outputs are analysed by the Long-range Dependencies Transformer (LDT) that captures the relations between the representations of the different image regions. For training the proposed network, we propose the Stream Drop augmentation. This randomly discards some network streams during training to prevent the network from over-attending on specific forensic signals. Our proposed approach is employed both for feature and score level fusion, combining RGB information with two and five forensic signals, respectively. We demonstrate its effectiveness on five popular datasets, compared with several handcrafted and deep learning state-of-the-art approaches. Moreover, we demonstrate its robustness against common perturbations and recent neural filters. An overview of the approach is presented in Fig. 1.
@@ 3.5 Training process
Feature Fusion | SPAN [30] | 21.4 | 83.3 | 72.5 | 98.1 | 42.5 | 86.5 | 0.4 | 56.2 | 0.2 | 52.8 | 20.8 | 75.0 | 7.5 | 61.5 | 23.6 | 73.3 |
@@
OMG-FuserF (Ours) | 84.5 | 97.2 | 86.1 | 97.2 | 63.1 | 91.4 | 34.2 | 79.1 | 72.2 | 92.1 | 53.9 | 88.3 | 77.0 | 95.4 | 67.3 | 91.5 | |
Score Fus. | DST-Fusion [21] | 75.3 | 94.8 | 85.4 | 93.5 | 48.8 | 79.2 | 14.7 | 56.9 | 20.4 | 50.7 | 38.6 | 80.9 | 24.9 | 72.8 | 44.0 | 75.6 |
@@
Feature Fusion | PSCC-Net [47] | 54.0 | 86.9 | 26.6 | 80.8 | 29.7 | 65.7 | 63.6 | 62.0 | 52.9 | 77.8 | 45.4 | 74.6 |
@@
OMG-FuserF (Ours) | 90.7 | 96.5 | 99.2 | 99.9 | 78.8 | 83.5 | 71.4 | 78.2 | 72.0 | 81.7 | 82.4 | 88.0 | |
Score Fus. | DST-Fusion [21] | 80.3 | 93.83 | 92.3 | 97.9 | 69.0 | 78.4 | 44.4 | 75.1 | 66.0 | 78.5 | 70.4 | 84.7 |
@@ 4.1 Experimental setup
OMG-Fuser provides a network for combining both the score-level outputs of multiple image forensics algorithms as well as multiple low-level image forensic signals. Thus, we experiment with two different variants of our architecture: (i) one that incorporates five recent forgery localization algorithms [30, 74, 12, 42, 25] (OMG-FuserS), and (ii) one that directly fuses two recently proposed learnable forensic cues, namely the DCT stream [42] and the Noiseprint++[25] at the feature-level (OMG-FuserF). We train both variants on 40k forged samples and 32k authentic samples compiled from the datasets provided in [41, 54, 13], while we evaluate them on CASIAv1+ [12], Columbia [29], Coverage [71], NIST16 [24], OpenForensics [43], CocoGlide [25] and DID [73] datasets. We employ six recent competitor methods with publicly available implementations for the feature fusion category [30, 47, 74, 12, 42, 25]. Regarding score-level fusion, we reimplement two popular methods that support the fusion of multiple signals [7, 21], along with a baseline average fusion approach, using the same input signals as OMG-FuserS. We use the SAM [36] pretrained model for instance segmentation. DINOv2 [59] model is used as the vision backbone, which is fine-tuned on our OMG-FuserF variant, but kept frozen on OMG-FuserS. Moreover, we employ the F1-score with a threshold of as the main indicator of performance when deployed in the wild and the AUC as an auxiliary threshold-agnostic metric. More information regarding implementation details, datasets, and algorithms, along with localization results in terms of the F1 metric using the best threshold per image, is provided in the supplementary material.
@@ 4.2 Comparison with the state-of-the-art
Image forgery localization evaluation results are presented in Tab. 1. The reported F1 and AUC metrics are computed on the pixel-level. We see that our implementations outperform all state-of-the-art methods in score- and feature-level fusion. Moreover, in the case of score-level fusion, we show that previous approaches, either based on some statistical frameworks or learning approaches without semantic information, perform poorly on the recent deep learning-based forensic signals, even worse than averaging the signals. This effectively shows that our network is capable of effectively fusing multiple forensic signals by exploiting the object-level information of the image, and at the same time, it is capable of handling low-level forensic traces. The high performance in feature-level fusion enables future works on new forensic signals to use our network to decrease the burden of developing purpose-built fusion architectures.
@@
OMG-FuserS | 70.4 | 93.5 | 83.2 | 89.5 | |
Signals | w/o SPAN | 69.2 | 91.5 | 82.3 | 89.1 |
@@
w/o Stream Drop | 66.5 | 90.5 | 81.1 | 87.3 | |
OMG-FuserF | 67.3 | 91.5 | 82.4 | 88.0 | |
Signals | w/o Noiseprint++ | 60.7 | 85.2 | 72.9 | 83.4 |
w/o DCT | 62.0 | 85.3 | 73.4 | 83.7 | |
w/o RGB | 64.3 | 82.8 | 77.6 | 84.7 |
@@
Image forgery detection evaluation results are presented in Tab. 2. Following [25], for methods that do not output an image-level forgery detection score, we compute it as the max of the corresponding forgery localization mask. Also, we omit datasets that contain only forged samples. Both OMG-Fuser implementations outperform all state-of-the-art methods by a significant margin.
@@ 4.3 Ablation Study
In order to better understand the contribution of each of the major components of our architecture, we performed several ablation studies in two directions. First, we evaluated the contribution of each of the fused forensic signals by training our network from scratch and removing the inputs of the network one at a time. Similarly, we proceeded by evaluating the contribution of each of the key components of our architecture by: replacing the OGA with the standard unmasked self-attention [68], replacing the TFT with an average pooling layer, removing the LDT and Stream Drop components. To better understand the impact of each component on signal fusion, we performed the later study on the OMG-FuserS, which fuses the output of several forensic algorithms. The results of the ablation study are presented in Tab. 3. Removing a single input signal affects the performance of the network adversely. Thus, the network is capable of learning to exploit even minor additional information introduced by new input streams. Moreover, we see that removing any key component of the architecture considerably impacts performance, highlighting that they are of crucial importance for effective signal fusion.
@@ 4.4 Robustness against Perturbations and Filters
4.5 Stream Expansion
The design of the TFT allows the expansion of an already trained model with additional streams. In Tab. 4, we report the overall performance across all the evaluation datasets when expanding our architecture in three different ways. Starting from a variant of OMG-FuserS trained without the TruFor stream, we expand it i) by training only the additional stream and keeping the rest of the network frozen, ii) by fine-tuning the network for 15% of the initial training epochs, and iii) by training the network from scratch. We see that in the first two cases, the network achieves a performance very close to third by requiring only a portion of the computational resources. This suggests that our model has learned to fuse various forensic signals effectively; therefore, it can be expanded with new signals without training from scratch.
@@ 5 Conclusions
Acknowledgments: This work was supported by the Horizon Europe vera.ai project (grant no.101070093), Junior Star GACR (grant no. GM 21-28830M), and EuroHPC for providing access to MeluXina supercomputer (grant no. EHPC-DEV-2023D03-008).
@@
Appendix A Experimental Setup
A.1 Image Forensics Algorithms
In our evaluation, we consider only methods with publicly available code in order to fairly evaluate them under the same setup. We employ as inputs to our score-level fusion network, OMG-FuserS, the SPAN [30], MVSS-Net++ [12], CATNetv2 [42], TruFor [25] and IFOSN [74] algorithms, due to their competitive performance and their complementarity with respect to the type of artifacts they detect, i.e., artifacts in the RGB domain, edge artifacts, compression artifacts, noise artifacts, and robustness against online sharing operations, respectively. Moreover, we utilize the Noiseprint++ [25] and the DCT-domain stream [42] as inputs to our feature-level fusion network, OMG-FuserF, two learnable forensic signals that captures noise-related and compression-related artifacts, respectively. For all the aforementioned methods, we utilize the official code implementations provided by the original authors.
A.2 Datasets
To train our OMG-Fuser models, we utilize data from three publicly available datasets. We randomly sample 25k forged and 25k authentic images from the synthetic dataset used by CAT-Net [41] due to the big variability of its samples regarding compression qualities and depicted topics. We enrich this set with another 10k inpainted images sampled from the DEFACTO [54] dataset. Also, we include all images from the CASIAv2 [13] dataset into our training set to compensate for low-resolution and low-quality images. We utilize 90% of these data for the actual training of the model and the rest 10% for validation purposes. Despite some datasets consisting of more samples, we observed in our experiments that a further increase in the amount of training data did not yield any significant performance increase. Thus, our architecture requires significantly less training data than the previous state-of-the-art ones [25, 42]. Instead, as highlighted by [25], introducing more variability into the training data was more beneficial. For evaluation, following [30, 47, 25], we have selected five popular datasets, namely the CASIAv1+ [12], COLUMBIA [29], COVERAGE [71], NIST16 [24], OpenForensics [43] datasets, including common cases of image forgery. In contrast to previous works [30, 47], we benchmark our methods on the entire evaluation dataset without making any further assumptions about the type of forgery or utilizing subsets of them. Furthermore, to take into account deep-learning based inpainting operations, we further employed the recently introduced CocoGlide [25] dataset and the deep-learning based inpaintings of the DiverseInpaintingDataset (DID) [73]. A summary of all the datasets used in our research is presented in Tab. 5.
@@ A.2 Datasets
A.3 Implementation Details
@@
Feature Fusion | PSCC-Net [47] | 83.4 | 86.7 | 69.3 | 50.2 | 51.3 | 84.1 | 58.5 | 70.7 |
@@
OMG-FuserF (Ours) | 92.0 | 94.6 | 88.3 | 82.1 | 82.0 | 87.7 | 78.9 | 80.1 | |
Score Fus. | DST-Fusion [21] | 89.3 | 91.8 | 76.6 | 56.0 | 33.3 | 86.5 | 64.8 | 77.4 |
@@
OMG-FuserS (Ours) | 92.2 | 96.6 | 89.3 | 83.7 | 85.2 | 86.2 | 86.1 | 86.8 |
For comparison with other score-level fusion approaches, we employ the OW-Fusion [7], a deep learning-based fusion approach, and implement it with the same input signals used on OMG-FuserS. Furthermore, in order to take into account the previous statistical fusion frameworks, we reimplement a DST-based fusion framework [21, 61] in Python, using again the same inputs with our score-level fusion implementation. Finally, we use the average of all input signals as a baseline approach.
Appendix B Additional Experiments
@@
OMG-FuserF (Ours) | 32.3 ms | |
Score Fusion | SPAN [30] | 1.34s |
@@
Appendix C Explainability
To better understand which parts of the input signals contribute the most to the fused tokens (eq. 5), we employ the Grad-CAM [65] method. In particular, we compute the gradients of the fused tokens with respect to the different inputs to the TFT in (eq. 4) in order to isolate the token fusion process and determine from which tokens the information propagates to the next stages. We compute the gradients based on the output of the TFT, using the squared -norm of the . In these experiments, we employ the OMG-FuserS variant of our architecture due to the easier interpretation of the input signals. The explainability maps for three samples are presented in Fig. 7. Our architecture has learned to attend to the correctly estimated regions in the input signals based on the ground truth while ignoring the regions of the input signals containing erroneous predictions. Also, our network focuses on the signals that better capture the forged and the authentic regions separately, e.g. on MVSSNet++ for the detection of authentic regions, while exploiting information from the image to resolve ambiguity in the input signals.
Appendix D Discussion on Research Ethics